Hi, here's the question.. Using the zero method, solve the equation graphically. 9x+7+5x+1=13x+17
Basically, what I think they're asking is 'solve for x.'
Thank in advance.

Answer by vertciel(183)

Hello there,
Yes, I agree with you in that you are asked to solve for x. However, I do not understand why you are asked to use the zero method for an equation of one degree (no exponents).
So I just did it algebraically:
9x + 7 + 5x +1 = 13x+17
13x - 9x - 5x = 7 + 1 - 17
-x = -9

See if you can find what x is, I found -x for you.